A car is traveling at a constant speed of 18 m/s on a highway. At the instant this car passes an entrance ramp, a second car enters the highway from the ramp. The second car starts from rest and has a constant acceleration. What acceleration must it maintain, so that the two cars meet for the first time at the next exit, which is 1.9 km away?
Answer #1

d = v01 t + 1/2 a1 t2 = v01 t

t = d / v01 -------------(1)

d = v02 t + 1/2 a2 t2 = 1/2 a2 t2

substitute (1) in above equation

d = 1/2 a2 (d / v01)2 = 1/2 a2 [d2 / v012]

a2 = 2 v012 / d

= (2 * 182) / (1.9 * 103)

acceleration the second car maintain = 0.341 m/s2